The relationship shown in the table is a direct variation.
x = 5, y = 15; x = 6, y = 18; x = 7, y = 21; x = 8, y = 24
Which equation best represents this relationship?
f y = 4x - 5
g y = x + 10
h y = 3x
j y = (1/3)x

Answer by Edwin McCravy(20054)

Plug in x=5 in them all to see if we get y = 15.

 f y = 4x - 5 = 4(5)-5 = 20-5 = 15.  That works:
 g y = x + 10 = 5+10 = 15.  That works.
 h y = 3x = 3(5) = 15.  That works,
 j y = (1/3)x = (1/3)(5) = 5/3.  That doesn't work.  So we eliminate j


Plug in x=6 in the remaining ones to see if we get y = 18.

 f y = 4x - 5 = 4(6)-5 = 24-5 = 19.  That doesn't work. So we eliminate f
 g y = x + 10 = 6+10 = 16.  That doesn't work.  So we eliminate g
 h y = 3x = 3(6) = 18.  That works,  So that's the correct choice.

Edwin